# Word problems and embedding-obstructions in cellular automata groups on groups

### v1

@@ 1.1 Complexity-theoretic results
Let $G$ be a f.g. group with at least stretched exponential growth, and let $X\subset A^{G}$ be any SFT where some finite-support point has free orbit. Then $\mathrm{Aut}(X)$ has a finitely-generated subgroup with PSPACE-hard word problem.
@@
For example $X=A^{G}$ satisfies the assumptions on $X$. We give more generally a sufficient geometric condition for such an upper bound in terms of “polynomial splitting schemes” for balls. We leave it as an open question what groups admit such schemes.
@@
The group $\mathbb{Z}_{p}\wr\mathbb{Z}$ itself has word problem decidable in polynomial time. We believe the analog of Theorem E can be proved similarly on solvable Baumslag-Solitar groups $\mathrm{BS}(1,n)=\mathbb{Z}[\frac{1}{n}]\rtimes\mathbb{Z}$, although we do not show this.
The lamplighter group and Baumslag-Solitar groups are somehow fundamental examples among solvable groups. First, of course they are metabelian, i.e. the derived length is minimal among nonabelian groups. Furthermore by a result of Groves [18], every f.g. virtually solvable group is either virtually nilpotent, or there exists a section (quotient of a subgroup) of the form $\mathbb{Z}_{p}\wr\mathbb{Z}$, $\mathrm{BS}(1,n)$ (a Baumslag-Solitar group) or $\mathbb{Z}^{d}\rtimes_{A}\mathbb{Z}$ for a matrix $A$ with an eigenvalue $>1$.
Of course the complexity of the word problem can likely become harder when we take a quotient of the spatially acting group, [^1] In the case of free groups $F_{k}$ we show PSPACE word problems for CA groups, which by a standard conjecture does not contain co-NEXPTIME, which in turn is the complexity of the word problem for some quotients of $F_{k}$ as we show. but still the result of Groves provides some motivation to look at groups $\mathbb{Z}^{d}\rtimes_{A}\mathbb{Z}$, and suggests that there might be some hope in attacking larger classes of solvable groups. We mention by analog that the domino problem, whether a given SFT is nonempty, is know to be undecidable on $\mathbb{Z}_{p}\wr\mathbb{Z}$, $\mathrm{BS}(1,n)$ [4] and (not virtually cyclic) polycyclic groups including $\mathbb{Z}^{d}\rtimes_{A}\mathbb{Z}$ [21].
@@
Of course, there are many general constructions that allow making the word problem of an f.g. group as difficult as one likes while preserving its decidability. A strong result in this direction is [22].
@@ 1.2 Non-embeddability results
Besides complexity theoretic results, we obtain some non-embeddability results.
@@
The following is a question (or infinite family of questions) of Hochman [19]:
@@
Does $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}^{D}})$ embed in $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}^{d}})$ for $d\neq D$?
@@
As noted in [2], the groups $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ and $\mathrm{Aut}(A^{\mathbb{Z}^{D}})$ are non-isomorphic for $d\neq D$, since their centers are the respective shift groups $\mathbb{Z}^{D}$ and $\mathbb{Z}^{d}$. This is a result of Hochman [19] (generalizing Ryan’s theorem [27], and further generalized by Barbieri, Carrasco-Vargas and Rivera-Burgos in [2]).
@@
Denote by $F_{2}$ the free group on two generators. The following is a question of Barbieri, Carrasco-Vargas and Rivera-Burgos [2]:
@@
Theorem G.
Let $G$ be a f.g. group containing a free submonoid on two generators, and let $A$ be any alphabet with sufficiently many divisors. Then $\mathrm{Aut}(A^{G})$ does not embed into $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any alphabet $B$ and $d\in\mathbb{N}$.
@@ Acknowledgements
We thank Ilkka Törmä for working through the proof of Theorem 5.1 in full detail. Of course, any mistakes are mine. The connection between the Gap Conjecture and a PSPACE frontier originated from the study of another computational model with Ilkka Törmä (apart from the construction of markers). We thank Sebastián Barbieri for sharing Question 1.2, and Scott Schmieding for several inspiring discussions. We thank Laurent Bartholdi for helping streamline the surface group argument (Lemma 9.2).
@@ 2 Definitions, conventions and some standard observations
$d(g,h)=\min\{n\;|\;h=gs_{1}s_{2}\ldots s_{n}\mbox{ for some }s_{i}\in S\},$
The commutator of group elements $a,b$ is $[a,b]=a^{-1}b{-1}ab$; $[G,G]$ is the group generated by commutators $[a,b]$ where $a,b\in G$. Also define longer commutators by $[a_{1},a_{2},\ldots,a_{k}]=[a_{1},[a_{2},\ldots,a_{k}]]$. A group $G$ is nilpotent if the lower central series $G,G_{1}=[G,G],G_{2}=[G,G_{1}],G_{3}=[G,G_{2}],\ldots$ reaches the trivial group in finitely many steps. A group $G$ is solvable if the derived series $G,G^{1}=[G,G],G^{2}=[G^{1},G^{1}],\ldots$ reaches the trivial group in finitely many steps.
Two groups $G,H$ are isomorphic if there is a bijection between them that preserves the multiplication relation, and we write this as $G\cong H$. The groups $G,H$ are commensurable if there exist subgroups $L\leq G,L^{\prime}\leq H$ of finite index such that $L\cong L^{\prime}$. The index of $H\leq G$ is the cardinality of $G/H=\{gH\;|\;g\in G\}$.
A group $G$ is virtually P (where P is any group, property of groups or family of groups) if $G$ has a subgroup of finite index which is P. It is residually P if for all $g\in G$ there is a quotient in $P$ where $g$ maps nontrivially. A group $G$ is amenable if for all $\epsilon>0$ and $S\subset G$ finite, there exists finite $F\subset G$ such that $\frac{FS}{F}<1+\epsilon$. A group $G$ is left-orderable if there exists a total order $<$ on $G$ such that $a<b\implies ga<gb$ for all $a,b$.
@@ 4 Groups with free submonoids
We show that $\mathcal{G}$ is a subgroup of $\mathrm{PAut}=\mathrm{PAut}(\mathbb{Z}_{7},B,\{0,1\})$, as required in the statement. This is only nontrivial for $\phi_{\pi,s,g}$. To show it is in $\mathcal{G}$, we will use the $\{0,1,2\}$-component of the state to “move” a permutation from one cell to another. The $\{0,1,2\}$-component acts here as a borrowed bit in the terminology [37] (or perhaps borrowed trit would be more accurate); this means that while it is used in the construction, its initial value is irrelevant and this value is returned at the end.
We now give the detailed proof of $\phi_{\pi,s,g}\in\mathcal{G}$. Write $\pi=[\pi_{0},\pi_{1}]$ where neither $\pi_{i}$ has $s$ in its support. We define four automorphisms by describing their effect on the symbol $x_{c}$ of the input $x\in A^{G}$, for $c\in G$. Define an automorphism $\theta_{b}$ that, if the $\{0,1,2\}$-component of the alphabet at $x_{c}$ is $b\in\{0,1\}$, permutes the first component of $x_{c}$ by $\pi_{i}$. Clearly this is in $\mathrm{PAut}$. Then define an automorphism $\theta$ that permutes the $\mathbb{Z}_{3}$-component of $x_{c}$ by the rotation (0 1 2)012(0\;1\;2)( 0 1 2 ) if the $\mathbb{Z}_{7}$-component of $x_{cg}$ is $s$. This is conjugate to a symbol permutation by a partial shift, so again in $\mathrm{PAut}$. Finally, let $\theta^{\prime}$ be the automorphism that permutes the $\mathbb{Z}_{3}$-component of $x_{c}$ by the rotation (0 1 2)012(0\;1\;2)( 0 1 2 ) in any case. Note that
@@
Suppose then that does not apply the permutation. There are several reasons why this might not happen:
  • either we have that for some $\ell\in[0,2^{i}]$ and some $h\in\{a,b\}^{\ell}$, the node $ch$ does not have counter value $j2^{i}+\ell$; or
@@ 5 Unconditional nonembedding results
The following lemma is the basis of our nonembedding results. It is a simple333pun intended way to express that the size of quotients needed in the definition of residual finiteness of $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ has an upper bound in terms of the radii of automorphisms.
@@
Let $A$ be a finite alphabet, let $\mathcal{G}=\langle S\rangle\leq\mathrm{Aut}(A^{\mathbb{Z}^{d}})$, where $S$ is a finite generating set defining the word norm of $\mathcal{G}$. Then there exists $C\in\mathbb{N}$ such that whenever $H\leq\mathcal{G}$ is a simple group containing a nontrivial element of norm at most $n$, we have $|H|\leq\mathrm{Sym}(C^{n^{d}})$.
In particular, $|H|\leq\mathrm{Sym}(C^{n^{d}})$ implies that $H$ has cardinality at most $|C|^{n^{d}}!$.
@@
Let $G$ be a f.g. group containing a free submonoid on two generators, and let $A$ be any alphabet with sufficiently many divisors. Then $\mathrm{Aut}(A^{G})$ has an f.g. subgroup that does not embed into $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any alphabet $B$ and $d\in\mathbb{N}$.
@@
Let first $A$ have many prime divisors. In the case of a general alphabet, the marker argument is similar as in the previous theorem.

### v2

@@ 1.1 Complexity-theoretic results
Let $G$ be a f.g. group with at least stretched exponential growth, and let $X\subset A^{G}$ be any SFT where some finite-support point has free shift orbit. Then $\mathrm{Aut}(X)$ has a finitely-generated subgroup with PSPACE-hard word problem.
@@
For example $X=A^{G}$ satisfies the assumptions on $X$. We give more generally a sufficient geometric condition for such an upper bound in terms of “polynomial splitting schemes” for balls. We leave it as an open question what groups admit such schemes. One interesting question is whether they exist for small cancellation groups or hyperbolic groups with one-dimensional boundary.
@@
The group $\mathbb{Z}_{2}\wr\mathbb{Z}$ itself has word problem decidable in polynomial time. We believe the analog of Theorem E can be proved similarly on solvable Baumslag-Solitar groups $\mathrm{BS}(1,n)=\mathbb{Z}[\frac{1}{n}]\rtimes\mathbb{Z}$, although we do not show this.
We mention by analog that the domino problem, whether a given SFT is nonempty, is known to be undecidable on $\mathbb{Z}_{2}\wr\mathbb{Z}$, $\mathrm{BS}(1,n)$ [4].
@@
Some interesting cases to look at would be polycyclic groups (that are not virtually cyclic), hyperbolic groups and $F_{2}\times F_{2}$.
The fact that we show the word problems to be this difficult for “natural” groups may be of interest in itself. There are certainly many general constructions that allow making the word problem of an f.g. group as difficult as one likes while preserving its decidability. A strong result in this direction is [22].
@@ 1.2 Non-embeddability results
Besides complexity theoretic results, we obtain some non-embeddability results from the same construction.
@@
The following was asked by Hochman in the conference “Current Trends in Dynamical Systems and the Mathematical Legacy of Rufus Bowen” organized in Vancouver in 2017. (A similar question, but allowing an arbitrary SFT in place of $\{0,1\}^{\mathbb{Z}^{d}}$ was asked in [19].)
@@
Does $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}^{D}})$ embed in $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}^{d}})$ for $d<D$?
@@
As noted in [2], the groups $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ and $\mathrm{Aut}(A^{\mathbb{Z}^{D}})$ are non-isomorphic for $d\neq D$, since their centers are the respective shift groups $\mathbb{Z}^{D}$ and $\mathbb{Z}^{d}$. This is a result of Hochman [19] (generalizing Ryan’s theorem [27], and further generalized by Barbieri, Carrasco-Vargas and Rivera-Burgos in [2]). However, it seems impossible to obtain any results about embeddability using this kind of method.
Hochman asked in personal communication whether Theorem F is also true modulo the centers. Indeed, this is the case, by a minor modification of the proof.
Theorem G.
Suppose that $D>3d+1$, $\mathcal{G}_{d}=\mathrm{Aut}(A^{\mathbb{Z}^{d}}),\mathcal{G}_{D}=\mathrm{Aut}(B^{\mathbb{Z}^{D}})$ for any nontrivial alphabets $A,B$. Then $\mathcal{G}_{D}$ has a finitely-generated subgroup such that $\mathcal{G}_{D}/K$ does not embed in $\mathcal{G}_{d}/K^{\prime}$ for any central subgroups $K\leq\mathcal{G}_{D},K^{\prime}\leq\mathcal{G}_{d}$.
In particular, the groups are not isomorphic modulo their centers.
Next, denote by $F_{2}$ the free group on two generators. The following is a question of Barbieri, Carrasco-Vargas and Rivera-Burgos [2]:
@@
Theorem H.
Let $G$ be a f.g. group containing a free submonoid on two generators, and let $A$ be any nontrivial alphabet. Then $\mathrm{Aut}(A^{G})$ does not embed into $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any alphabet $B$ and $d\in\mathbb{N}$.
@@ Acknowledgements
We thank Ilkka Törmä for working through the proof of Theorem 4.1 in full detail. Of course, any mistakes are mine. The connection between the Gap Conjecture and a PSPACE frontier originated from the study of another computational model with Ilkka Törmä (apart from the construction of markers). We thank Sebastián Barbieri for sharing Question 1.2, and Scott Schmieding for several inspiring discussions. We thank Laurent Bartholdi for helping streamline the surface group argument (Lemma 8.2), and other useful comments. We thank Michael Hochman for asking about Theorem G.
@@ 2.1 Assorted definitions
2.2 Groups
$d(g,h)=\min\{n\;|\;h=gs_{1}s_{2}\ldots s_{n}\mbox{ for some }s_{i}\in S\},$
The commutator of group elements $a,b$ is $[a,b]=a^{-1}b{-1}ab$; $[G,G]$ is the group generated by commutators $[a,b]$ where $a,b\in G$. Also define longer commutators by $[a_{1},a_{2},\ldots,a_{k}]=[a_{1},[a_{2},\ldots,a_{k}]]$. A group is abelian if any two elements $g,h$ commute. A group $G$ is nilpotent if the lower central series $G,G_{1}=[G,G],G_{2}=[G,G_{1}],G_{3}=[G,G_{2}],\ldots$ reaches the trivial group in finitely many steps. A group $G$ is solvable if the derived series $G,G^{1}=[G,G],G^{2}=[G^{1},G^{1}],\ldots$ reaches the trivial group in finitely many steps. The center of a group $G$ is the set of elements $g\in G$ that commute with all other elements.
Two groups $G,H$ are isomorphic if there is a bijection between them that preserves the multiplication relation, and we write this as $G\cong H$. The groups $G,H$ are commensurable if there exist subgroups $L\leq G,L^{\prime}\leq H$ of finite index such that $L\cong L^{\prime}$. The index $[G:H]$ ofa subgroup $H\leq G$ is the cardinality of $G/H=\{gH\;|\;g\in G\}$.
A group $G$ is virtually P (where P is any group, property of groups or family of groups) if $G$ has a subgroup of finite index which is P. It is residually P if for all $g\in G$ there is a quotient in $P$ where $g$ maps nontrivially. A group $G$ is amenable if for all $\epsilon>0$ and $S\subset G$ finite, there exists finite $F\subset G$ such that $\frac{FS}{F}<1+\epsilon$. A group $G$ is left-orderable if there exists a total order $<$ on $G$ such that $a<b\implies ga<gb$ for all $a,b,g\in G$.
2.3 Word problems
2.4 Symbolic dynamics
@@ 3 Groups with free submonoids
We show that $\mathcal{G}$ is a subgroup of $\mathrm{PAut}=\mathrm{PAut}(\mathbb{Z}_{7},B,\{0,1\})$, as required in the statement. This is only nontrivial for $\phi_{\pi,s,g}$. To show it is in $\mathrm{PAut}$, we will use the $\{0,1,2\}$-component of the state to “move” a permutation from one cell to another. The $\{0,1,2\}$-component acts here as a borrowed bit in the terminology [37] (or perhaps borrowed trit would be more accurate); this means that while it is used in the construction, its initial value is irrelevant and this value is returned at the end.
We now give the detailed proof of $\phi_{\pi,s,g}\in\mathrm{PAut}$. Write $\pi=[\pi_{0},\pi_{1}]$ where neither $\pi_{i}$ has $s$ in its support. We define four automorphisms by describing their effect on the symbol $x_{c}$ of the input $x\in A^{G}$, for $c\in G$. Define an automorphism $\theta_{b}$ that, if the $\{0,1,2\}$-component of the alphabet at $x_{c}$ is $b\in\{0,1\}$, permutes the first component of $x_{c}$ by $\pi_{i}$. Clearly this is in $\mathrm{PAut}$. Then define an automorphism $\theta$ that permutes the $\mathbb{Z}_{3}$-component of $x_{c}$ by the rotation (0 1 2)012(0\;1\;2)( 0 1 2 ) if the $\mathbb{Z}_{7}$-component of $x_{cg}$ is $s$. This is conjugate to a symbol permutation by a partial shift, so again in $\mathrm{PAut}$. Finally, let $\theta^{\prime}$ be the automorphism that permutes the $\mathbb{Z}_{3}$-component of $x_{c}$ by the rotation (0 1 2)012(0\;1\;2)( 0 1 2 ) in any case. Note that
@@
Suppose then that $\psi_{j,i,s,\pi}$ does not apply the permutation. There are several reasons why this might not happen:
  • we may have that for some $\ell\in[0,2^{i}]$ and some $h\in\{a,b\}^{\ell}$, the node $ch$ does not have counter value $j2^{i}+\ell$; or
@@ 4 Unconditional nonembedding results
The following lemma is the basis of our nonembedding results. It is a simple way to express that the size of quotients needed in the definition of residual finiteness of $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ has an upper bound in terms of the radii of automorphisms.
@@
Let $A$ be a finite alphabet, let $\mathcal{G}=\langle S\rangle\leq\mathrm{Aut}(A^{\mathbb{Z}^{d}})$, where $S$ is a finite generating set defining the word norm of $\mathcal{G}$. Then there exists $C\in\mathbb{N}$ such that whenever $H\leq\mathcal{G}$ is a simple group containing a nontrivial element of norm at most $n$, we have $H\leq\mathrm{Sym}(C^{n^{d}})$.
In particular, $H\leq\mathrm{Sym}(C^{n^{d}})$ implies that $H$ has cardinality at most $|C|^{n^{d}}!$.
@@
Let $G$ be a f.g. group containing a free submonoid on two generators, and let $A$ be any alphabet. Then $\mathrm{Aut}(A^{G})$ has an f.g. subgroup that does not embed into $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any alphabet $B$ and $d\in\mathbb{N}$.
@@
Assume $A$ has many prime divisors. In the case of a general alphabet, the marker argument is similar as in the previous theorem.
@@
4.1 Dropping the centers
By completely analogous arguments, one can prove variants of the results where the center is quotiented out. We omit a technical version of Lemma 4.1, and explain the modifications needed directly.
The following is due to Hochman [19]. The one-dimensional result is due to Ryan [27].
Lemma 4.2.
The center of $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ is the group of shifts.
Theorem 4.3.
Suppose that $D>3d+1$, $\mathcal{G}_{d}=\mathrm{Aut}(A^{\mathbb{Z}^{d}}),\mathcal{G}_{D}=\mathrm{Aut}(B^{\mathbb{Z}^{D}})$ for any nontrivial alphabets $A,B$. Then $\mathcal{G}_{D}$ has a finitely-generated subgroup such that $\mathcal{G}_{D}/K$ does not embed in $\mathcal{G}_{d}/K^{\prime}$ for any central subgroups $K\leq\mathcal{G}_{D},K^{\prime}\leq\mathcal{G}_{d}$.
Proof.
The proof is analogous to Theorem 4.1. Let $\mathcal{G}=\langle S\rangle$ be the group used in that proof. Then in $\mathcal{G}_{D}/K$ we take the the subgroup $\mathcal{G}/K$.
A we showed, there is a subgroup $H$ of $\mathcal{G}_{D}$ which is nonabelian simple of order $27^{\Theta(n^{D-1})}!/$ 2, and contains an element of word norm $n$. Now, $H$ must have trivial intersection with $K$, since otherwise $K\cap H$ is a nontrivial simple subgroup of $H$. Thus by the second isomorphism theorem, $HK/K\cong H$ as a subgroup of $\mathcal{G}/K$, and we conclude that $\mathcal{G}/K$ still has a simple subgroup of order $27^{\Theta(n^{D-1})}!/$ 2 and with an element of word norm $n$.
Next we observe that the set of groups embeddable in some $\mathcal{G}_{d}/K^{\prime}$ (over all alphabets $B$) does not decrease if we assume $K^{\prime}$ is just the subgroup $\langle e_{1},\ldots,e_{d^{\prime}}\rangle$ of $\mathbb{Z}^{d}$ for some $d^{\prime}\leq d$. Namely, we observe first (e.g. by linear algebra) that there is a subgroup $L$ such that $G=K^{\prime}\oplus L$ is of finite index in $\mathbb{Z}^{d}$. Any subgroup of $\mathcal{G}_{d}/K^{\prime}=\mathrm{Aut}(B^{\mathbb{Z}^{d}},\mathbb{Z}^{d})/K^{\prime}$ is certainly a subgroup of $\mathrm{Aut}(B^{\mathbb{Z}^{d}},G)/K^{\prime}$. Then we recall that the $G$-action $(B^{\mathbb{Z}^{d}},G)$ itself is topologically conjugate to a full shift on symbols $B^{[\mathbb{Z}^{d}:G]}$. By replacing the alphabet $B$ by this set and observing that $G\cong\mathbb{Z}^{d}$ (by the fundamental theorem of abelian groups), we may replace the original group $\mathbb{Z}^{d}$ by $G$. Since $K^{\prime}\times L$ is now an internal direct product in $\mathbb{Z}^{d}$, by the fundamental theorem of abelian groups $K^{\prime}\cong\mathbb{Z}^{d^{\prime}}$ and $K^{\prime}\cong\mathbb{Z}^{d^{\prime\prime}}$ for some $d^{\prime}+d^{\prime\prime}=d$, and thus by changing the basis we may assume $K^{\prime}$ is just the subgroup $\langle e_{1},\ldots,e_{d^{\prime}}\rangle$.
Assume now that $K^{\prime}$ is of this form, and let $L=\langle e_{d^{\prime}+1},\ldots,e_{d^{\prime\prime}}\rangle$. Consider a finitely-generated subgroup $H^{\prime}$ of $\mathcal{G}_{d}/K^{\prime}$. We may take representatives for the generators $\mathcal{S}^{\prime}\subset\mathcal{G}_{d}$, and suppose $r$ is a bound on their radii. Suppose that some nontrivial element $h\in H^{\prime}$ of this group has word norm $n$ and belongs to a nonabelian simple group of order $27^{\Theta(n^{D-1})}!/$ 2. Let $h^{\prime}$ be the corresponding representative in $\mathcal{G}_{d}$ (of the same word norm). Since $h$ is nontrivial, $h^{\prime}$ is not contained in $K^{\prime}$. Note that the radius of $h^{\prime}$ is at most $rn$.
We now claim that we can find a point with periods at most $10rn$ in each direction, such that $h^{\prime}$ does not act trivially on the $K^{\prime}$-orbits of periodic points. We observe that $\mathcal{G}_{d}/K^{\prime}$ has a well-defined action on the $K^{\prime}$ orbits of such points, and thus so does $H^{\prime}$. Since $H^{\prime}$ is simple, as in the proof of Lemma 4.1 we see that $H^{\prime}$ must act faithfully on the set of $K^{\prime}$-orbits, and we conclude that again $|H^{\prime}|\leq C^{n^{d}}!$ for some constant $C$ (which is only a function of $r$ and the alphabet. Note here that the constant $10$ also disappears into $C$, and that the number of $K^{\prime}$-orbits is smaller than the number of periodic points, which only gives a better upper bound for $H$.
There are two cases to consider for this claim. Suppose $B\ni 0,1$. First, suppose $h^{\prime}$ is a shift by an element not in $K^{\prime}$, then it is a shift by a vector of radius at most $rn$ from outside $K^{\prime}$. We can easily find point with periods $(2rn+1)$ where $h^{\prime}$ it does not act by a $K^{\prime}$-shift, for example we can pick the point $x\in\{0,1\}^{\mathbb{Z}^{d}}$ with periods $2(rn+1)\mathbb{Z}^{d}$, where in the fundamental domain $[0,2rn]^{d}$ we have $x_{\vec{v}}=1\iff\vec{v}\in K$.
Suppose then that $h^{\prime}$ is not a shift map. Let first $y\in\{0,1\}^{\mathbb{Z}^{d}}$ be the characteristic function of $\{\vec{0}\}$, i.e. $y_{\vec{v}}=1\iff\vec{v}=\vec{0}$. Then we have two subcases. First, if $h^{\prime}(y)\notin K^{\prime}\cdot y$, then it is easy to see that we can pick the point $x\in\{0,1\}^{\mathbb{Z}^{d}}$ with periods $2(rn+1)\mathbb{Z}^{d}$, where in the fundamental domain $[0,2rn]^{d}$ we have $x_{\vec{v}}=1\iff\vec{v}=\vec{0}$, and then $x$ is not mapped into its $K^{\prime}$-shift orbit.
Suppose then $h^{\prime}(y)\in K^{\prime}\cdot y$, so $h^{\prime}(y)=\sigma_{\vec{v}}(y)$ for some $\vec{v}\in B_{rn}\cap K^{\prime}$. Since $h^{\prime}$ is not a shift by an element of $K^{\prime}$, we can find a pattern $P\in B^{B_{rn}}$ such that $h^{\prime}(z)\neq\sigma_{\vec{v}}(z)$ when $z\in[P]$. Positioning $z|B_{rn}$ and a central pattern of $y$ far from each other in the same $10rn\mathbb{Z}^{d}$-periodic point $x$, we obtain a point which $h^{\prime}$ does not shift by an element of $K^{\prime}$.
This concludes the proof, since as explained above, this gives a bound of the form $|H|\leq C^{n^{d}}!$ as needed in the theorem. ∎